### v1

# Odd coloring graphs with linear neighborhood complexity

Abstract
We prove that any class of bipartite graphs with linear neighborhood complexity has bounded odd chromatic number. As a result, if $\mathcal{G}$ is the class of all circle graphs, or if $\mathcal{G}$ is any class with bounded twin-width, bounded merge-width, or a forbidden vertex-minor, then $\mathcal{G}$ is $\chi_{odd}$-bounded.
1 Introduction
The odd chromatic number $\chi_{odd}(G)$ of a graph $G$ is the minimum number of colors needed to properly color $G$ so that for each non-isolated vertex $v$ of $G$, there exists a color which appears on an odd number of vertices in its open neighborhood $N(v)$. This parameter [^1] Note that in the literature there exists another parameter called odd chromatic number, see e.g. [2]. was recently introduced by Petruševski and Škrekovski [37]. Among the main problems and results concerning this parameter is the conjecture of [37] that all planar graphs have odd chromatic number at most $5$, where $8$ is the best current upper bound [36]. An open problem from [10] states that the odd chromatic number is bounded by the maximum degree plus a constant, where current best results can be found in [11]. From a sparsity point of view, classes of bounded expansion have bounded odd chromatic number. Indeed, they even have bounded proper conflict free-chromatic number [27, 33] and bounded strong odd chromatic number [38]; see the end of this section for the relevant definitions.
In the current paper, we prove that any bipartite graph of linear neighborhood complexity has bounded odd chromatic number. There are many examples of graphs with linear neighborhood complexity. In particular, classes of graphs with good model-theoretic properties often have linear neighborhood complexity; see, for example, [3, 4, 20, 40]. The neighborhood complexity of a graph $G$, denoted by $\eta_{G}(m)$, is the maximum over all $m$-element subsets $A\subseteq V(G)$ of |$|\{N(v)\cap A:v\in V(G)\}|$. Thus, the neighborhood complexity counts the number of different neighborhoods that can occur within an $m$-vertex set.
Theorem 1.1.
For any integer $r$, there exists an integer $k=k(r)$ so that if $G$ is a bipartite graph with neighborhood complexity $\eta_{G}(m)\leqslant r\cdot m$ for every positive $m$, then $\chi_{odd}(G)\leqslant k$.
The main ingredient in our proof of Theorem 1.1 is Haussler’s Shallow Packing Lemma [26].
We obtain the following corollary of Theorem 1.1 using a technique from [29] to reduce to the case of bipartite graphs. A graph class is $\chi_{odd}$-bounded if there exists a function $f$ so that any graph in the class with clique number $\omega$ has odd chromatic number at most $f(\omega)$.
Corollary 1.2.
If $\mathcal{G}$ is any class of graphs with bounded twin-width, bounded merge-width, or a forbidden vertex-minor, then $\mathcal{G}$ is $\chi_{odd}$-bounded.
Our proof of Corollary 1.2 uses the fact that classes with bounded twin-width [5, 9, 39], bounded merge-width [3], or a forbidden vertex-minor [13] are $\chi$-bounded. (The chromatic number of a graph $G$ is denoted by $\chi(G)$, and a class of graphs is $\chi$-bounded if there exists a function $f$ so that any graph $G$ in the class with clique number $\omega$ has $\chi(G)\leqslant f(\omega)$.)
Next, we use a connection between the Matroid Growth Rate Theorem [22] and neighborhood complexity in order to obtain the following corollary of Theorem 1.1. This connection was recently observed by the authors [14], who used it in order to find the unavoidable cosimple $\textsf{GF}(q)$-representable matroids of large girth.
Corollary 1.3.
For any integers $t$ and $q$, there exists an integer $f(t,q)$ such that for any matroid $M$ without $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ as a minor and for any basis $B$ of $M$, the fundamental graph $\mathcal{F}(M,B)$ has odd chromatic number at most $f(t,q)$.
An interesting common special case of Corollary 1.2 and Corollary 1.3 is that of (bipartite) circle graphs. A graph is a circle graph if there exists a collection of chords on a circle so that the vertices are in bijection with the chords, and two vertices are adjacent if and only if their chords have non-empty intersection. Bouchet [8] characterized circle graphs by their forbidden vertex-minors. Additionally, circle graphs with bounded clique number have bounded twin-width [28]. On the other hand, bipartite circle graphs are the fundamental graphs of planar graphs [16]. We also note that there are several direct proofs that circle graphs are $\chi$-bounded; see [12, 15, 25, 30].
Hence, the above results settle [29, Question 3.9], which asked whether circle graphs are $\chi_{odd}$-bounded. Using a technique completely independent of the previous results, we further improve the bound on the odd chromatic number of bipartite circle graphs to 98. This is a more explicit bound, even if it is far from the lower bound; see 5.3.
Corollary 1.4.
Every bipartite circle graph has odd chromatic number at most 98.
We obtain this result as a corollary of the more general Theorem 5.2, which can be seen as a “one-sided” odd-coloring result for the fundamental graph of a graphic matroid. We believe that Theorem 5.2 is interesting on its own and propose a possible strengthening; see 5.4.
[…]
Outline
In Section 2, we introduce Haussler’s Shallow Packing Lemma and some preliminaries about matroids. We prove Theorem 1.1 in Section 3. In Section 4, we prove the two corollaries of Theorem 1.1 (Corollary 1.2 and Corollary 1.3). Finally, in Section 5, we use a different technique to find a concrete bound on the odd chromatic number of bipartite circle graphs and propose a couple of open problems.
2 Preliminaries
We write $K_{t}$ for the complete graph with $t$ vertices, $M(G)$ for the cycle matroid of a graph $G$, and $M^{*}$ for the dual of a matroid $M$. We write $U_{2,q}$ for the $q$-element line, and, more generally, $U_{t,q}$ for the uniform matroid with $q$ elements and rank $t$. That is, $U_{t,q}$ is the $q$-element matroid where the circuits are the sets of size $t+1$. Given a graph $G$ and a vertex $v$, we write $N_{G}(v)$ (or just $N(v)$ when the graph is clear from the context) for the open neighborhood of $v$. That is, we do not consider $v$ to be in its neighborhood $N(v)$.
Haussler’s Shallow Packing Lemma
Let $\mathcal{F}$ be a collection of subsets of a finite ground set $V$. We always consider $\mathcal{F}$ as a subset of $2^{V}$; we do not allow multisets. The shatter function of $\mathcal{F}$, denoted by $\pi_{\mathcal{F}}(m)$, is the maximum size of $\mathcal{F}$ when restricted to any $m$ elements in $V$. That is, $\pi_{\mathcal{F}}(m)$ is the maximum, over all $m$-element subsets $W\subseteq V$, of the number of equivalence classes of the relationship $\sim_{W}$ on $\mathcal{F}$ where two elements $F,F^{\prime}\in\mathcal{F}$ satisfy $F\sim_{W}F^{\prime}$ if $F\cap W=F^{\prime}\cap W$. For a positive integer $\delta$, we say that two sets $F,F^{\prime}\in\mathcal{F}$ are $\delta$-separated if their symmetric difference $F\Delta F^{\prime}$ has size at least $\delta$ (that is, there are at least $\delta$ elements in $V$ which are in one of $F,F^{\prime}$ but not the other). We say that $\mathcal{F}$ is $\delta$-separated if any pair of distinct elements in $\mathcal{F}$ are $\delta$-separated.
The following version of Haussler’s Shallow Packing Lemma from [26] is stated as [19, Lemma 2.2], for instance. We just require the case that $d=1$.
Lemma 2.1 ([26]).
For any number $r\geqslant 1$, there exists an integer $c=c(r)$ so that for every positive integer $\delta$, if $\mathcal{F}$ is a collection of subsets of a finite ground set $V$ so that $\mathcal{F}$ is $\delta$-separated and $\pi_{\mathcal{F}}(m)\leqslant rm$ for every positive integer $m$, then $|\mathcal{F}|\leqslant c|V|/\delta$.
Fundamental Graphs
Let $M$ be a matroid. Given a basis $B$ of $M$, the fundamental circuit of an element $e\in E(M)\setminus B$ with respect to $B$ is the unique circuit of $M$ which is contained in $B\cup\{e\}$. The fundamental graph $\mathcal{F}(M,B)$ is the bipartite graph with sides $B$ and $E(M)\setminus B$ where each element $e\in E(M)\setminus B$ is adjacent to the other elements in its fundamental circuit. Given a graph $G$, the bases of the cycle matroid of $G$ are the edge sets of spanning forests; a spanning forest is a maximal acyclic subgraph of $G$. If $T$ is a spanning forest of $G$, then the fundamental graph of $G$ with respect to $T$ is the graph $\mathcal{F}(M(G),E(T))$. For convenience, we denote this graph by $\mathcal{F}(G,T)$.
Bouchet [7] introduced fundamental graphs for binary matroids and showed how they behave with respect to taking matroid minors. However, since Bouchet used different (more general) terminology, we refer the reader to Oum [35] for a nice introduction to fundamental graphs of binary matroids. For general matroids, see for instance [24].
Lemma 2.2 ([24, Proposition 3.1(ii)]).
For any matroid $M$ and basis $B$ of $M$, we have $\mathcal{F}(M,B)=\mathcal{F}(M^{*},E(M)\setminus B)$.
For a matroid $M$ and element $e\in E(M)$, we write $M\setminus e$ for the matroid obtained from $M$ by deleting $e$ and $M/e$ for the matroid obtained from $M$ by contracting $e$. For a graph $G$ and a vertex $v\in V(G)$, we write $G-v$ for the graph obtained from $G$ by deleting $v$. Consider deleting a vertex from the non-basis side of a fundamental graph; this corresponds to just deleting that element from the matroid. Thus, we obtain the following lemma, using Lemma 2.2 and the fact that for any matroid $M$ and element $e$, we have $M^{*}\setminus e=(M/e)^{*}$.
Lemma 2.3.
For any matroid $M$, basis $B$ of $M$, and element $e\in E(M)$,
$\mathcal{F}(M,B)-e=\begin{cases}\mathcal{F}(M\setminus e,B)&\textrm{if }e\in E(M)\setminus B\cr\mathcal{F}(M/e,B\setminus\{e\})&\textrm{if}e\in B.\end{cases}$
In Section 4, we use the Growth Rate Theorem of Geelen and Whittle [22] to prove that the fundamental graphs of certain matroids have linear neighborhood complexity (see Lemma 4.1). The Growth Rate Theorem of Geelen and Whittle [22] says the following. (We remark that there is also a more general Growth Rate Theorem of Geelen, Kung, and Whittle [21].)
Theorem 2.4 ([22]).
For any integers $t$ and $q$, there exists an integer $g(t,q)$ such that any simple rank-$n$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor has at most $g(t,q)\cdot n$ elements.
Note that graphic matroids exclude the line $U_{2,4}$ as a minor. We remark that Nelson, Norin, and Omana [34] have recently improved the bound on $g(t,q)$ in Theorem 2.4 to a single exponential function.
3 Odd coloring and neighborhood complexity
In this section we prove Theorem 1.1, which says that bipartite graphs of linear neighborhood complexity have bounded odd chromatic number.
First we need to prove a lemma which says that graphs of linear neighborhood complexity contain vertices which have almost the same neighborhood. Let $G$ be a graph, and let $\delta$ be an integer. We say that two vertices $u,v\in V(G)$ are $\delta$-near-twins if $|N(u)\Delta N(v)|\leqslant\delta$, where we write $\Delta$ for the symmetric difference. Likewise, we say that two vertices are twins if they have the exact same neighborhood.
We now observe that linear neighborhood complexity is enough to obtain a pair of $\mathcal{O}(1)$-near-twins within the desired side of a bipartite graph (under some mild additional assumptions). In fact, we show that we can force the near-twins to be within any linearly-sized subset of that side of the bipartite graph.
Lemma 3.1.
For any integers $r$ and $\ell$, there exists an integer $\delta=\delta(r,\ell)$ so that if $G$ is a bipartite graph with a bipartition $(X,Y)$ so that $X$ does not contain any pair of twins and $\eta_{G}(m)\leqslant r\cdot m$ for every positive integer $m$, then for every $Y^{\prime}\subseteq Y$ with $|Y^{\prime}|\geqslant\max\{2,|Y|/\ell\}$, there exist distinct vertices $u,v\in Y^{\prime}$ which are $\delta$-near-twins in $G$.
Proof.
Let $c=c(r)$ be the constant from Haussler’s Shallow Packing Lemma, which is stated as Lemma 2.1. Then set $\delta\coloneqq r\ell c+1$.
Going for a contradiction, suppose there are no two vertices in $Y^{\prime}$ which are $\delta$-near-twins in $G$. Let =$\mathcal{F}=\{N(v):v\in Y^{\prime}\}$, so that $\mathcal{F}$ is a collection of subsets of $X$. Since $Y^{\prime}$ does not contain any pair of twins, $|\mathcal{F}|=|Y^{\prime}|$. Moreover, since $Y^{\prime}$ does not contain any pair of $\delta$-near-twins, $\mathcal{F}$ is $\delta$-separated. Finally, by the assumption about the neighborhood complexity of $G$, we have that $\pi_{\mathcal{F}}(m)\leqslant rm$ for every positive $m$. So $|Y^{\prime}|\leqslant c|X|/\delta$ by Lemma 2.1. However, since $X$ contains no twins, we also have
$|X|\leqslant|\{N(v)\cap Y:v\in V(G)\}|\leqslant r|Y|\leqslant r\ell|Y^{\prime}|\leqslant r\ell c|X|/\delta.$
Thus $\delta\leqslant r\ell c$, however we chose $\delta=r\ell c+1$, a contradiction. ∎
Now we are ready to prove Theorem 1.1. We restate the theorem below for convenience.
See 1.1
Proof.
Let $(X,Y)$ be a bipartition of $G$. We claim that it suffices to show how to color $Y$ so that every non-isolated vertex in $X$ has a color that appears an odd number of times in its neighborhood. Then, if $\varphi_{Y}:Y\rightarrow\mathbb{N}$ is the coloring of $Y$, and $\varphi_{X}:X\rightarrow\mathbb{N}$ is the coloring of $X$ (which exists by symmetry between $X$ and $Y$), we can take the final coloring to be
$\phi(v)=\begin{cases}(1,\varphi_{X}(v))&\textrm{if }v\in X\cr(2,\varphi_{Y}(v))&\textrm{if }v\in Y.\end{cases}$
Taking this product coloring ensures that $\phi$ is proper.
Let $\delta$ be the function from Lemma 3.1, and set $k\coloneqq\max\left(2r+1,\delta(r,2r+1)+1\right)$. We now show that such a coloring $\varphi_{Y}$ exists using at most $k$ colors. Going for a contradiction, suppose otherwise. Let $G$ be a graph with a bipartition $(X,Y)$ which forms a counterexample. Choose $G,X,Y$ so that $|V(G)|$ is minimum. First, we prove two claims.
Claim 1.
There is no pair of twins in $X$.
Proof.
If $X$ contains two vertices $u$ and $v$ which are twins, then by minimality there exists a coloring $\varphi_{Y}:Y\rightarrow\mathbb{N}$ so that every non-isolated vertex in $X\setminus\{v\}$ has a color that appears an odd number of times in its neighborhood. This same coloring works for $G$. ∎
Our goal is to apply Lemma 3.1 to find a pair of near-twins in $Y$. Then we will use the same color for those two vertices, and apply the fact that $G$ is a minimum counterexample. However, we will have to avoid pairs of vertices $u,v\in Y$ so that there exists a degree-$2$ vertex in $X$ whose neighbors are $u$ and $v$. (Such a vertex forces us to use different colors for $u$ and $v$.) The next claim gives us a large subset of $Y$ to work with.
Claim 2.
There exists a set $Y^{\prime}\subseteq Y$ of size at least $|Y|/(2r+1)$ so that no degree-$2$ vertex in $X$ has both its neighbors in $Y^{\prime}$.
Proof.
Consider an auxiliary graph $H$ with vertex-set $Y$ where two vertices $u,v\in Y$ are adjacent if there exists a vertex in $X$ whose neighbors are precisely $u$ and $v$. We claim that $H$ is $2r$-degenerate. This will complete the proof since then we can find an independent set of size at least $|Y|/(2r+1)$ by taking a greedy coloring and choosing the largest color class. So, consider an induced subgraph $H^{\prime}$ of $H$. The number of edges of $H^{\prime}$ is at most the number of different neighborhoods $G$ has within $V(H^{\prime})$. So $|E(H^{\prime})|\leqslant r\cdot|V(H^{\prime})|$ by our bound on the neighborhood complexity of $G$. Thus $H$ is $2r$-degenerate, as desired. ∎
Let $Y^{\prime}\subseteq Y$ be as in Claim 2. Notice that $|Y^{\prime}|\geqslant 2$ since otherwise $|Y|\leqslant 2r+1$ and we could color every vertex in $Y$ with a different color. Now, notice that $G,X,Y,Y^{\prime}$ satisfy the conditions of Lemma 3.1 with $\ell\coloneqq 2r+1$. (Recall that $X$ contains no pair of twins by Claim 1.) Thus, there exist distinct vertices $u,v\in Y^{\prime}$ which are $(k-1)$-near-twins. By the minimality of $G$, there exists a coloring $\varphi:Y\setminus\{u,v\}\rightarrow\mathbb{N}$ with at most $k$ colors so that every vertex in $X$ with at least one neighbor in $Y\setminus\{u,v\}$ has a color that appears an odd number of times in its neighborhood. By the choice of $Y^{\prime}$, this is actually every vertex in $X$ except for possibly some vertices of degree 00 or $1$. We do not need to worry about the vertices in $X$ of degree 00 or $1$ as they automatically satisfy the condition.
Let us say that a color $i$ is bad if there exists a vertex $x\in N(u)\Delta N(v)$ so that $i$ is the unique color that appears an odd number of times in $N(x)\cap(Y\setminus\{u,v\})$. Thus, there are at most $|N(u)\Delta N(v)|\leqslant k-1$ bad colors. It follows that there exists a coloring $\varphi_{Y}$ of $Y$ which is obtained from $\varphi$ by coloring $u$ and $v$ the same color, which is not bad. It follows that for every non-isolated vertex $x\in X$ there is a color that appears an odd number of times in $N(x)$. (Note that if $x$ is adjacent to both of $u$ and $v$ or neither of $u$ and $v$, then the number of times each color appears in $N(x)$ has the same parity as in $N(x)\setminus\{u,v\}$.) ∎
4 Applications
In this section we prove the two corollaries of Theorem 1.1 (Corollary 1.2 and Corollary 1.3).
We begin by using the Growth Rate Theorem (stated as Theorem 2.4) to control the neighborhood complexity of certain fundamental graphs. We note that we could do a “one-sided” version of the following lemma. However this version, which is symmetric under taking duals, is more convenient for our purposes. Note that Corollary 1.3 immediately follows from Theorem 1.1 and the following lemma.
Lemma 4.1.
For any integers $t$ and $q$, there exists an integer $r=r(t,q)$ so that if $M$ is a matroid with no $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ minor, and $B$ is a basis of $M$, then the neighborhood complexity of $\mathcal{F}(M,B)$ is at most $r\cdot m$ for every positive $m$.
Proof.
Let $g$ denote the function from Theorem 2.4; so any simple rank-$n$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor has at most $g(t,q)\cdot n$ elements. Let $M$ be a matroid with no $U_{2,q+2}$, $U_{q,q+2}$, $M(K_{t})$, or $M(K_{t})^{*}$ minor, and let $B$ be a basis of $M$.
Since $\mathcal{F}(M,B)$ is bipartite, and a matroid and its dual have the same fundamental graph by Lemma 2.2, it suffices to consider the number of different neighborhoods within a set $A\subseteq B$. Let $F\subseteq E(M)\setminus B$ be obtained by selecting one vertex $v$ with $N(v)\cap A=A^{\prime}$ for each subset $A^{\prime}\subseteq A$ for which $|A^{\prime}|\geqslant 2$, and such a vertex $v$ exists. Thus |$|\{N(v)\cap A:v\in V(\mathcal{F}(M,B))\}|\leqslant|F|+|A|+1$ because $A$ has $|A|+1$ subsets with at most one element.
Now, let $M^{\prime}$ be the matroid obtained from $M$ by contracting all elements in $B\setminus A$ and deleting all elements in $(E(M)\setminus B)\setminus F$. By Lemma 2.3, the subgraph of $\mathcal{F}(M,B)$ induced on $A\cup F$ is equal to $\mathcal{F}(M^{\prime},A)$. Note that $M^{\prime}$ is a rank-$|A|$ matroid with no $U_{2,q+2}$ or $M(K_{t})$ minor. We claim that $M^{\prime}$ is simple as well. This will complete the proof since then the Growth Rate Theorem implies that $|F|+|A|=|E(M^{\prime})|\leqslant g(t,q)\cdot|A|$.
First of all, $M^{\prime}$ has no loops since $A$ is a basis and every vertex in $F$ has at least one neighbor in $A$. Next, $M^{\prime}$ has no parallel pair with one element in $A$ and the other in $F$ since every element in $F$ has at least two neighbors in $A$. Finally, $M^{\prime}$ has no parallel pair with both elements in $F$ since no two vertices in $F$ have the same neighborhood in $A$. So $M^{\prime}$ is simple, which completes the proof. ∎
Recall that a class of graphs is $\chi_{odd}$-bounded if there exists a function $f$ so that any graph in the class with clique number $\omega$ has odd chromatic number at most $f(\omega)$. Similarly, a class is $\chi$-bounded if there exists a function $g$ so that any graph in the class with clique number $\omega$ has chromatic number at most $g(\omega)$. We denote the chromatic number of a graph $G$ by $\chi(G)$. We say that a class is hereditary if it is closed under taking induced subgraphs.
The following lemma from [29] implies that in order to understand whether a hereditary $\chi$-bounded class is also $\chi_{odd}$-bounded, it suffices to consider the bipartite graphs in the class. We state a slightly weaker but simpler version of the lemma.
Lemma 4.2 ([29, Lemma 1.3]).
If every induced bipartite subgraph of a graph $G$ has odd chromatic number at most $t$, then $\chi_{odd}(G)\leqslant t^{2}\chi(G)$.
Now we may prove Corollary 1.2. We omit the definitions of twin-width, merge-width, and vertex-minors; instead we refer the reader to [6][18], and [13], respectively. The corollary is restated below for convenience.
See 1.2
Proof.
We may assume that $\mathcal{G}$ is hereditary because taking induced subgraphs cannot increase the twin-width or merge-width, or create a new vertex-minor. Recall that classes with bounded twin-width [5, 9, 39], classes with bounded merge-width [3], and classes with a forbidden vertex-minor [13] are known to be $\chi$-bounded. Thus, by Lemma 4.2 and Theorem 1.1, it suffices to prove that every bipartite graph $H\in\mathcal{G}$ has linear neighborhood complexity. For classes of bounded twin-width, this was proven by Bonnet, Foucaud, Lehtilä, and Parreau [4]. For classes with bounded merge-width, this was recently proven by Bonamy and Geniet [3].
For classes with a forbidden vertex-minor, we observe that the $1$-subdivision of $K_{t}$ contains every $t$-vertex graph as a vertex-minor, and is itself the fundamental graph of $K_{t+1}$ with respect to the $(t+1)$-vertex star. Since every pivot-minor is also a vertex-minor, it follows that any class of bipartite graphs with a forbidden vertex-minor also forbids the $1$-subdivision of $K_{t}$ as a pivot-minor (for some $t$). Finally, due to the connection between pivot-minors of bipartite graphs and minors of binary matroids [35], it follows from Lemma 4.1 that bipartite graphs with a forbidden vertex-minor have linear neighborhood complexity, as desired. ∎
5 Odd coloring and bipartite circle graphs
In the following we consider an asymmetric situation, in which we want to color only $E\setminus B$ in $\mathcal{F}(M,B)$ such that every non-isolated vertex from $B$ has some color occurring an odd number of times in its neighborhood. In fact, we only consider this setting in graphic matroids of graphs with no $K_{t}$-minor. That is, we give another proof that for any $t$ there is a constant $k=k(t)$ so that given any graph $G$ with no $K_{t}$-minor and any spanning forest $T$ of $G$, we can color the non-forest edges with $k$ colors so that every non-bridge edge $e$ has a color which appears an odd number of times in its neighborhood in $\mathcal{F}(G,T)$.
Given a graph $G$ and a spanning forest $T$ of $G$, the fundamental cut of an edge $f\in E(T)$ is the set $C^{*}(T,f)$ which contains all edges in $E(G)$ which join the vertex-sets of the two components of $T-f$ which are not components of $T$. Note that $f$ itself is included in $C^{*}(T,f)$. Indeed, $C^{*}(T,f)\setminus\{f\}$ is the neighborhood of $f$ in $\mathcal{F}(G,T)$.
Lemma 5.1.
For any graph $G$, any spanning forest $T$ of $G$, and any edge $f_{0}$ of $T$, there exists a coloring $\varphi:C^{*}(T,f_{0})\setminus\{f_{0}\}\rightarrow\{1,\ldots,9\}$ so that for each tree-edge $f\in E(T)$ with $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\neq\emptyset$ there exists a color $i\in\{1,\ldots,8\}$ which occurs an odd number of times on $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})$.
Proof.
Let $T_{1}$ and $T_{2}$ be the two components of $T-f_{0}$. We shall color the non-tree edges in the fundamental cut $C^{*}(T,f_{0})$ of $f_{0}$, i.e., the non-tree edges with one endpoint in $T_{1}$ and the other endpoint in $T_{2}$. Consider the tree $T_{1}^{\prime}=T_{1}+f_{0}$ to be rooted at the endpoint of $f_{0}$ in $T_{2}$. Observe that for any edge $f$ in $T_{1}^{\prime}$, a non-tree edge $e\in C^{*}(T,f_{0})$ lies in $C^{*}(T,f)$ if and only if the endpoint of $e$ in $T_{1}$ lies in the subtree of $T_{1}^{\prime}$ below $f$. It follows that the set family $\mathcal{S}_{1}=\{C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\colon f\in T_{1}^{\prime}\}$ is laminar, i.e., its members are either disjoint or in a containment relation. From [29, Lemma 3.8] we get that there is a coloring $\varphi_{1}:C^{*}(T,f_{0})\setminus\{f_{0}\}\to\{1,2,3\}$ such that every element of $\mathcal{S}_{1}$ has an odd number of elements colored $1$ or an odd number of elements colored $2$.
Symmetrically, there is a coloring $\varphi_{2}:C^{*}(T,f_{0})\setminus\{f_{0}\}\to\{1,2,3\}$ such that every element of $\mathcal{S}_{2}=\{C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\colon f\in T_{2}^{\prime}\}$ has an odd number of elements colored $1$ or an odd number of elements colored $2$. Setting $\varphi(e)=(\varphi_{1}(e),\varphi_{2}(e))$ for every $e\in C^{*}(T,f_{0})\setminus\{f_{0}\}$, we obtain the desired $9$-coloring. Indeed, if $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})$ has an odd number of elements colored $i$ in $\varphi_{1}$, then in $\varphi$ these elements are distributed into three color classes $(i,1)$, $(i,2)$, $(i,3)$, at least one of which is necessarily odd. Note that since by [29, Lemma 3.8] in both factors of the product coloring the color $3$ is not needed in order to witness having an odd color in the neighborhoods, we can designate color $(3,3)$ as color $9$ from the statement. ∎
For the next result, recall that the star-arboricity $\mathrm{sa}(G)$ of a graph $G$ is the smallest $\ell$, such that $E(G)$ can be decomposed into $\ell$ star forests, i.e., each component has diameter at most $2$. Note that the $\mathrm{sa}(G)$ is bounded by twice the degeneracy of $G$, see [1]. Hence, in particular for any integer $t$, there exists an integer $\ell(t)$ so that if $G$ is a graph with no $K_{t}$ minor, then $\mathrm{sa}(G)\leqslant\ell(t)$. Indeed, by the Kostochka–Thomason bound $\ell(t)\in O(t\sqrt{\log t})$, see [31, 32, 42, 41].
Theorem 5.2.
For any integer $t$, there exists an integer $k=k(t)$ so that if $G$ is a graph with no $K_{t}$ minor, and $T$ is a spanning forest of $G$, then there exists a coloring $\phi:E(G)\setminus E(T)\rightarrow\{1,\ldots,k\}$ so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,\ldots,k\}$ which occurs an odd number of times on $C^{*}(T,f)$. Indeed, $k(t)\leqslant 16\ell(t-1)+1$, with $\ell(\cdot)$ defined as above the theorem.
Proof.
Note that we may assume that $G$ is connected and $T$ is a spanning tree by considering the components separately. Further, we may assume that $G$ is $2$-edge-connected by contracting all bridges.
We root $T$ at an arbitrary vertex $r$. We shall construct a sequence $G_{0}\subseteq G_{1}\subseteq\cdots\subseteq G_{x}=G$ of subgraphs of $G$, starting with $G_{0}$ being just vertex $r$. For a non-tree edge $e$, the fundamental cycle $C(T,e)$ is the unique cycle in $T+e$. Note that for a tree-edge $f$ and a non-tree edge $e$ we have $f\in C(T,e)$ if and only if $e\in C^{*}(T,f)$. For $i>0$, having defined $G_{i-1}$, let
$G_{i}\coloneqq\bigcup\{C(T,e)\colon e\in E(G)\setminus E(T),C(T,e)\text{ contains a vertex of }G_{i-1}\}.$
Observe that for each $i\geqslant 0$ we have that $T_{i}=T\cap G_{i}$ is a subtree of $T$ containing the root $r$. For convenience, for each $i>0$ we set $A_{i}\coloneqq E(T_{i})-E(T_{i-1})$ as the set of tree-edges added in step $i$, and $B_{i}\coloneqq(E(G_{i})-E(G_{i-1}))-E(T)$ as the set of non-tree edges added in step $i$.
Claim 3.
If for a tree-edge $f\in A_{i}$ we have $C^{*}(T,f)\cap B_{j}\neq\emptyset$, then $i\leqslant j\leqslant i+1$.
Proof.
Consider any non-tree edge $e\in B_{j}$ that lies on the fundamental cut $C^{*}(T,f)$ of the tree-edge $f\in A_{i}$. Then, $f$ lies on the fundamental cycle $C(T,e)$, and it follows that $j\leqslant i+1$.
Conversely, since $e\in B_{j}$, the fundamental cycle $C(T,e)\subseteq G_{j}$ and thus $i\leqslant j$. ∎
Next, we color the edges in $B_{i}$ (which is non-empty only for $i>0$), i.e., the non-tree edges in $G_{i}$ that are not in $G_{i-1}$. To this end, consider $G_{i}$ with spanning tree $T_{i}$. Contract the subtree $T_{i-1}$ of $T_{i}$ to a single vertex $a$. Also, consider the components of $T_{i}$ after the removal of all vertices in $T_{i-1}$ and contract each such component to a single vertex. The resulting graph $H$ is a minor of $G_{i}$ with two vertices in $H-a$ being adjacent if and only if some non-tree edge in $B_{i}$ connects the corresponding subtrees of $T_{i}-V(T_{i-1})$.
Since $a$ is a universal vertex in $H$, the graph $H-a$ has no $K_{t-1}$ minor. Thus, as argued above the theorem there exists a universal constant $\ell=\ell(t-1)$ such that $E(H-a)$ can be decomposed into $\ell$ star forests $F_{1},\ldots,F_{\ell}$. (We may assume that each vertex of $H-a$ is the center of at least one star.) For each star $S$ in $F_{1},\ldots,F_{\ell}$, consider its center $v$ in $H-a$, the corresponding component $K$ of $T_{i}-V(T_{i-1})$, and the edge $t_{0}\in T_{i}$ that connects $K$ and $T_{i-1}$. Observe that $C^{*}(T,f_{0})\setminus\{t_{0}\}\subseteq B_{i}$. Let $B(S)\subseteq B_{i}$ be the subset of non-tree edges that correspond in $H$ to edges in the star $S$ or the edge between $a$ and $v$. By Lemma 5.1 there is a $9$-coloring $\varphi_{S}$ of $B(S)$ such that for every $f\in T$ the set $C^{*}(T,f)\cap(C^{*}(T,f_{0})\setminus\{f_{0}\})\cap B(S)$ is either empty or contains a color an odd number of times.
To color $B_{i}$, we use a new palette of $9$ colors for each of the $\ell$ star forests $F_{1},\ldots,F_{\ell}$, while reusing the same set $C_{x}$ of $9$ colors for all stars in the same star forest $F_{x}$. In fact, if $S_{1}$ and $S_{2}$ are two stars from the same forest $F_{x}$, then for every tree-edge $f\in A_{i}$ we have that $C^{*}(T,f)$ is either disjoint from $B(S_{1})$ or $B(S_{2})$ (or both). Hence, for every $f\in A_{i}$ we have that $C^{*}(T,f)\cap\left(\bigcup_{S\in F_{x}}B(S)\right)$ is either empty or contains a color in $C_{x}$ an odd number of times.
Finally, we use one set of $\ell$ palettes of $9$ colors to color $B_{i}$ for even $i$, and a second set of $\ell$ palettes of $9$ colors to color $B_{i}$ for odd $i$. By Claim 3, this results in a coloring $\phi$ with $2\cdot 9\cdot\ell$ colors that is odd for $\{C^{*}(T,f)\colon f\in T\}$. Using further, that by Lemma 5.1 in none of the palettes the color $9$ is used as the odd color in a neighborhood, we can merge all these, yielding a total of $12\cdot 8\cdot\ell+$ 1 colors. ∎
We now use Theorem 5.2 to prove Corollary 1.4, which is restated below.
See 1.4
Proof.
De Fraysseix [16] proved that every connected bipartite circle graph is a fundamental graph of a planar graph $G$ and a spanning tree $T$. So $G$ excludes $K_{5}$ as a minor, and the graphs excluding $K_{4}$ as a minor are exactly the graphs of treewidth at most $2$. Hence, by [17] they have star arboricity at most $3$. Thus, we can use Theorem 5.2 to color $E(G)\setminus E(T)$ with $16\cdot 3+1=49$ colors so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,2,\ldots,49\}$ which occurs an odd number of times on the fundamental cut of $f$.
Now, by planar duality $M(G^{*})\cong M(G)^{*}$, see e.g., [43, Chapter 6.3, Corollary 1]. Thus, by Lemma 2.2 the fundamental graph of $G$ and $T$ is isomorphic to the fundamental of graph $G^{*}$ with spanning forest $T^{*}=E(G)\setminus E(T)$. Applying Theorem 5.2 to this graph, we can use a second set of size $49$ to color $E(G^{*})\setminus E(T^{*})=E(T)$ such that for each non-bridge edge $f\in E(T^{*})=E(G)\setminus E(T)$, there exists a color $i\in\{50,51,\ldots,98\}$ which occurs an odd number of times on the fundamental cut of $f$ in $G^{*}$ with respect to $T^{*}$. ∎
We have shown that bipartite circle graphs have odd chromatic number at most 98. However, the best lower bound we know of is $4$, attained by the cycle of length $4$. It would be interesting to obtain tighter bounds.
Problem 5.3.
Does every bipartite circle graph have odd chromatic number at most $4$?
We think that Theorem 5.2 is interesting in its own right and in fact we are not aware of any family of graphs that forces an unbounded number of colors when coloring non-tree edges as in the theorem. In other words, we wonder if the following holds.
Problem 5.4.
Does there exist a constant $k$ so that if $G$ is a graph and $T$ is a spanning forest of $G$, then there exists a coloring $\phi:E(G)\setminus E(T)\rightarrow\{1,\ldots,k\}$ so that for each non-bridge edge $f\in E(T)$, there exists a color $i\in\{1,\ldots,k\}$ which occurs an odd number of times on the fundamental cut $C^{*}(T,f)$?
Acknowledgments
This work was completed at the 12th Annual Workshop on Geometry and Graphs held at the Bellairs Research Institute in February 2025. We are grateful to the organizers and participants for providing a excellent research environment. We would also like to thank Szymon Toruńczyk for helpful comments and Juan Pablo Peña for discussions on circle graphs.